OpenStudy (anonymous):

The surface area of an oil spill gets 105% larger every day, represented by the function s(x) = (1.05)x − 1. On the first day, it covered an area of 31 square meters. Which function would be used to find the area of the oil spill on the 15th day?

OpenStudy (anonymous):

@thushananth01

OpenStudy (anonymous):

I believe s(x) = (1.05)x - 15

OpenStudy (anonymous):

that's not one of the options :/ s(x) = (1.05)31 − 1 s(x) = 31(1.05)x − 1 s(x) = 1.05(31)x − 1 s(x) = (1.05)x − 31

OpenStudy (anonymous):

i think its d

OpenStudy (anonymous):

does anyone know for sure? I think its d too but idk

OpenStudy (anonymous):

I would say B, but I'm still trying to formulate a reason for it. Is it \(s(x)=1.05^x-1\) or \(s(x)=1.05x-1\)?

OpenStudy (anonymous):

ok just let me know when someone knows for sure because I want to get a good grade on this but theres another pn Every day, there are 4 times more likes on an internet video of a horse which is modeled by the function c(n) = (4)n − 1, where n is the number of days since the video posted. On the first day, there were 100 likes. What is the function that shows the number of likes each day?

OpenStudy (anonymous):

@thushananth01 anything for Every day, there are 4 times more likes on an internet video of a horse which is modeled by the function c(n) = (4)n − 1, where n is the number of days since the video posted. On the first day, there were 100 likes. What is the function that shows the number of likes each day?